OpenStudy (anonymous):

sin4theata=1 Can someone explain how to get this answer?

OpenStudy (vera_ewing):

\(\large\color{pink}{What~is~your~question?}\)

OpenStudy (anonymous):

\[\sin4\theta=1\] Solve each equation over the interval 0,2pi

OpenStudy (vera_ewing):

\(\large\color{pink}{What~do~you~think~the~answer~is?}\)

OpenStudy (anonymous):

N oidea

OpenStudy (anonymous):

Well, do you mind if I change it to sin(4x)=1?

OpenStudy (michele_laino):

hint: \[\sin \theta = 1,\quad \theta = \frac{\pi }{2} + 2k\pi \]

OpenStudy (michele_laino):

\[\sin x = 1,\quad x = \frac{\pi }{2} + 2k\pi \]

OpenStudy (michele_laino):

so we can write: \[4\theta = \frac{\pi }{2} + 2k\pi \] what is theta?

OpenStudy (anonymous):

4?

OpenStudy (jdoe0001):

any ideas?

OpenStudy (jdoe0001):

if you take \(\large sin^{-1}\) to both sides, first, what do you end up with?